# Surgical Aggregation: Federated Class-Heterogeneous Learning

### v5

@@ 3.1.2 Algorithm Overview
3.1.3 Convergence Analysis
Let us consider a multi-label classifier with initialized parameters $\mathbf{w}_{0,k}$ updated using gradient descent. We primarily focus this convergence analysis on the modified strategy for task-specific weights in our method.
Definition 3. We define task-specific weights in the task block $G$ as $(\mathbf{w}_{t,k}^{(G)})_{c}$, i.e., the weights corresponding to epoch $t\in[T]$ and client $k$ for the class $c\in C_{k}$.
We analyze the convergence of surgical aggregation under scenarios of class-heterogeneity described in Definition 2.
Let $(G\circ F)(\mathbf{w})$ be the global model, while $(G\circ F)_{k}(\mathbf{w})$ be the local model for client $k$, such that our distributed optimization problem is defined as:
$\min_{\mathbf{w}}\left\{(G\circ F)^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}(G\circ F)_{k}^{*}(\mathbf{w})\right\}$ (1)
@@ 3.1.3 Convergence Analysis
$(G\circ F)_{k}^{*}\triangleq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$ (2)
where $L$ is the loss function (e.g., binary cross-entropy) [18].
@@
$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (3)
@@
    Optimizing the global representation block $F$ by locally optimizing client representation blocks $F_{k}$, such that:
    $\min_{\mathbf{w}}\left\{F^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}^{*}(\mathbf{w})\right\}$ (4)
@@
    Optimizing the global task block $G$ by locally optimizing client task blocks $G_{k}$, such that:
    $\min_{\mathbf{w}}\left\{G^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}G_{k}^{*}(\mathbf{w})\right\}$ (5)
@@
Proof: Like conventional FL approaches, in our approach the representation block weights $\mathbf{w}_{k,t}^{(F)}$ are always shared with all clients at the end of each communication round, such that:
$\mathbf{w}_{t,k}^{(F)}=f(\mathbf{w}_{t}^{(F)})$ (6)
In this case, the convergence of $\mathbf{w}_{k,t}^{(F)}$ is well-established in prior literature [18, 17]. However, it is not clear whether class-heterogeneity affects the convergence of $\mathbf{w}_{k,t}^{(F)}$. The optimization of the global sub-objective $F^{*}$ is controlled by the optimization of the local objective $(G\circ F)^{*}_{k}$. Therefore, the optimization of $F^{*}$ is ensured if $F_{k}^{*}\triangleq(G\circ F)^{*}_{k}$ is optimized. While the optimization of $(G\circ F)^{*}_{k}$ is dependent on classes $C_{k}$, it is independent of the optimization of classes $c\in C\backslash C_{k}$ (Lemma 1). In other words, the optimization of $F^{*}$ is driven by the optimization of $F_{k}^{*}$, regardless of any class-heterogeneity.
Since the task block weights are selectively aggregated, we analyze the convergence of the global task block (i.e., $G^{*}$) with respect to each individual class $c\in C$.
@@
Proof: Let $K^{\prime}$ be the subset of clients in the network sharing class $c\in C$. In our approach, the task-specific weights $(\mathbf{w}_{k,t}^{(G)})_{c}$ for clients $k\in K^{\prime}$ are shared with the clients at the end of each communication round, such that:
$(\mathbf{w}_{k,t}^{(G)})_{c}=g((\mathbf{w}_{t}^{(G)})_{c})$ (7)
@@
Putting together Lemmas 3 and 4, we prove that the optimization of global objective $(G\circ F)^{*}$ with respect to class $c\in C$ is always ensured and is derived from previously established convergence analyses, regardless of class-heterogeneity in the network [18, 17]. Therefore, using Lemma 1, we can confer:
Theorem 1. The optimization of global objective $(G\circ F)^{*}$ is ensured if local objective $(G\circ F)_{k}^{*}$ is optimized for all classes $c\in C_{k}$ using representation block strategy $f$ and task block strategy $g$ and optimization is independent of any class-heterogeneity present in the network.

### v6

@@ 2.1.2 Overview of Surgical Aggregation
[…]
2.2 Datasets
[…]
Figure 4: Freqeuence of disease labels between NIH, CheXpert, and MIMIC datasets. The seven shared classes between the NIH and CheXpert datasets are bolded.
NIH Chest X-Ray 14: It consists of 14 disease labels with $n=112,120$ frontal CXRs from 30,805 patients (Wang et al., 2017). We randomly divide the dataset into training (70%, $n=78,075$), validation (10%, $n=11,079$), and testing (20%, $n=22,966$) splits with no patient leakage.
CheXpert: It consists of 13 disease labels (seven shared with NIH) with $n=224,316$ CXRs from 65,240 patients (Irvin et al., 2019). All lateral images are discarded to yield $n=191,027$ frontal CXRs. Uncertain labels are treated as negatives. We randomly divide the dataset into training (70%, $n=133,638$), validation (10%, $n=18,855$), and testing (20%, $n=38,534$) splits with no patient leakage.
MIMIC-CXR-JPG: It consists of $n=377,110$ CXRs from 65,379 patients and is used as our external, fully annotated test set to evaluate generalizability (Goldberger et al., 2000; Johnson et al., 2019). All lateral images are discarded to yield $n=243,324$ frontal CXRs. The CXR-LT expanded set of 26 labels provides ground-truth for all 20 classes present in NIH and CheXpert datasets (Holste et al., 2023).
Figure 5: Latent space representation of NIH, CheXpert, and MIMIC datasets using t-SNE (Van der Maaten and Hinton, 2008).
2.3 Baselines
We compare our method to two baselines that represent an upper bound of expected model performance, but do not yield a singular global model.
Individually Trained: Models trained individually on each dataset in a centralized setup using standard DL.
Personalized FL: PFL is a popular approach to FL, where each client learns a personalized local model in favor of generalized global models (Liang et al., 2020; Tan et al., 2022; Li et al., 2021). While clients share a feature extractor, they learn a personalized classification head for local classes.
We also compare our method to three existing methods for training models with class-heterogeneous datasets that result in a single global model.
Centralized: Multiple datasets are concatenated locally to train a model in a centralized setup using standard DL to achieve better generalizability (Cohen et al., 2020; Seyyed-Kalantari et al., 2021). All missing classes in the combined dataset are treated as negatives. For a client $k$ with local classes $C_{k}$, and data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$), $(y_{k,i})_{c}=0$ for all $c\not\in C_{k}$.
Vanilla FL: Conventional approach for FL where datasets are distributed across clients in the network (Chowdhury et al., 2022; Sheller et al., 2020). Missing classes in each dataset are treated as negatives.
FL with Partial Loss: A modification of the vanilla FL setup using a partial loss function that drops non-local classes when computing local losses (Jin and Ghahramani, 2002). For client $k$ with model parameters $\mathbf{w}_{t,k}$, local classes $C_{k}$, and private data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$), it is defined as:
$L(\mathbf{w}_{t,k},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w}_{t,k},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (4)
2.4 Experiments
2.4.1 Ablation Study
In this experiment, we evaluated the effect of 1) number of clients and 2) number of shared classes using simulated datasets. We sampled IID subsets of the NIH dataset to create simulated datasets for each client, such that there is no patient leakage and all 14 classes from the NIH dataset appear at least once in the network. Since each simulated dataset is an ”incomplete” snapshot of the NIH dataset, we evaluated the performance of surgical aggregation in relation to the NIH baseline model.
Effect of number of clients: We evaluated the effect of number of clients across seven iterations, each with an increasing number of clients $K\in\{2,3,4,5,6,8,10\}$. We randomly selected values of $K$, such that as $K$ increases, the sample size of the simulated dataset at each client $k$ decreases by factor of $1/K$.
Effect of number of shared classes: We evaluated the effect of number of shared classes as a proxy for class-heterogeneity in the network. Since the NIH dataset contains 14 classes, we randomly sampled seven possible iterations out of 14, each with an increasing number of shared classes $\{0,1,2,4,8,12,14\}$. The number of clients $K=4$ and underlying data at each client was constant.
All models were evaluated on the held-out NIH test set and the external MIMIC dataset. We used FedBN+ for all FL approaches (Kulkarni et al., 2023). The PFL baseline was not included because it does not train a global model.
2.4.2 Non-IID Experiments
We evaluated surgical aggregation using the NIH and CheXpert datasets across two sub-experiments:
[…]
Class-Heterogeneous: We considered a non-IID and class-heterogeneous setting, where all 20 classes from both datasets were included as ground-truth annotations. All models were evaluated on the NIH and CheXpert test sets using local classes (14 and 13 respectively) and the external MIMIC dataset using all 20 classes. We also measured the disparity in performance between the shared and unique classes for the NIH and CheXpert datasets using their respective test set. Furthermore, we measured the performance on the NIH and CheXpert classes when evaluating with the external MIMIC test set. This enabled us to compare our method with the individually trained and PFL baselines.
2.5 Implementation
Centralized Setup: We trained ImageNet-pretrained DenseNet121 models using binary cross-entropy (BCE) loss and a batch size of 64. Prior to training, the classification heads were ”warmed up” using transfer learning for 15 epochs with a learning rate of 5e-3, while keeping feature extractor weights frozen. Then, the model was trained with all weights unfrozen for 150 epochs with a learning rate of 5e-5.
Federated Setup: We simulated a FL setup with $T=150$ epochs and $E=1$ epochs before communication. Each client trained an ImageNet-pretrained DenseNet121 model using the methods described above. The classification heads of local models were similarly ”warmed up” prior to training. Global convergence was determined by the mean of local validation BCE loss for each client. We considered the three aggregation strategies: 1) FedAvg, the de-facto FL strategy (McMahan et al., 2017). 2) FedBN, a variation of FedAvg with local batch normalization (Li et al., 2021). As a PFL strategy, FedBN is only included for the PFL baseline model. 3) FedBN+, a modification of FedBN with pre-trained batch normalization statistics for global models (Kulkarni et al., 2023).
In order to primarily focus on class-heterogeneity, we used the same hyperparameters for all models trained. All CXRs were downsampled to $224\times 224$, normalized between 0 and 1, and scaled to ImageNet statistics. Random augmentations (rotation, flip, zoom, and contrast) were applied to input images during training. All models were trained and evaluated using TensorFlow (version 2.8.1) and CUDA (version 12.0) on four NVIDIA RTX A6000 GPUs.
Figure 6: Mean AUROC scores for ablation study evaluating (a) effect of number of clients, (b) effect of number of shared classes on the NIH (left) and MIMIC (right) test sets.
Figure 7: Average mean AUROC scores for the effect of number of clients on NIH and MIMIC test sets. (ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
Figure 8: Average mean AUROC scores for the effect of number of shared classes on NIH and MIMIC test sets. (ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
2.6 Metrics and Statistical Analyses
Model performance in a multi-label scenario is quantified by the mean of area under the receiver operating characteristic curve (AUROC) scores for each class in the test set. If a class in the test set is absent from the set of classes learned by a model, the mean AUROC score is not defined. All values are reported as Mean $\pm$ SD.
For the ablation study, performance between models is compared using paired t-tests between the mean AUROC scores to test for significance. For real-world experiments, performance between models is compared using paired t-tests across the granular class-level AUROC scores to test for significance. In all cases, statistical significance was defined as $p<0.05$.
3 Results
3.1 Ablation Study
Effect of number of clients: We observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p\leq 0.005$) (Figures 6(a) and 7). While FL with partial loss performs comparably to our method, the difference in average mean AUROCs ($\Delta=0.004$) is statistically significant ($p=0.004$). On the external MIMIC dataset, our method generalizes better with average mean AUROC $0.69\pm 0.01$, compared to all existing methods (all $p\leq 0.03$).
Effect of number of shared clients: We observe that surgical aggregation outperforms all existing methods with average mean AUROC $0.80\pm 0.01$ on the held-out NIH test set (all $p\leq 0.03$) (Figures 6(b) and 8). Similarly, the minor differences in average mean AUROCs between FL with partial loss and surgical aggregation are statistically significant ($\Delta=0.004$, $p=0.03$). On the external MIMIC dataset, our method generalizes better than all existing methods with average mean AUROC $0.69\pm 0.01$.
Finally, our results indicate that even with varying levels of class- and data-heterogeneity in the network, surgical aggregation provides the closest approximation to the performance of the NIH baseline model (mean AUROC $0.82$, mean $\Delta=0.017$), followed by FL with partial loss (mean $\Delta=0.021$).
Table 1: Mean AUROC scores on NIH, CheXpert, and MIMIC dataset for non-IID, non-class-heterogeneous experiment. Our method is compared with others using paired t-tests. Best performing method is bolded.
NIH CheXpert MIMIC
NIH $0.84\pm 0.05^{\text{ns}}$ $0.70\pm 0.07^{**}$ $0.73\pm 0.10^{**}$
CheXpert $0.81\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
PFL
FedAvg $0.83\pm 0.06^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
FedBN $0.83\pm 0.05^{*}$ $0.77\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
FedBN+ $0.83\pm 0.06^{**}$ $0.76\pm 0.07^{\text{ns}}$ $0.75\pm 0.08^{\text{ns}}$
Central $0.82\pm 0.06^{**}$ $0.77\pm 0.07^{\text{ns}}$ $0.76\pm 0.10^{\text{ns}}$
FL
FedAvg $0.81\pm 0.05^{***}$ $0.74\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
FedBN+ $0.84\pm 0.05^{*}$ $0.75\pm 0.08^{*}$ $0.75\pm 0.09^{\text{ns}}$
FL + PL
FedAvg $0.82\pm 0.05^{***}$ $0.75\pm 0.08^{*}$ $0.74\pm 0.10^{*}$
FedBN+ $0.83\pm 0.06^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.76\pm 0.09^{\text{ns}}$
FL + SA
FedAvg $0.84\pm 0.05^{*}$ $0.75\pm 0.07^{\text{ns}}$ $0.75\pm 0.09^{\text{ns}}$
PL: Partial Loss, SA: Surgical Aggregation
(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
3.2 Non-IID Experiments
[…]
On the NIH test set, our method performs comparably to the NIH baseline with mean AUROC $0.84\pm 0.06$ ($p=0.22$), while outperforming other baselines (all $p\leq 0.03$) and existing methods (all $p\leq 0.02$). On the CheXpert test set, our method performs comparably to the CheXpert baseline model with mean AUROC $0.77\pm 0.08$ ($p=0.51$). Our method performs on par with the PFL baselines (all $p>0.05$) and existing methods, while outperforming the NIH baseline ($p=0.002$)
Finally, on the external MIMIC dataset, we observe that surgical aggregation’s generalizability to unseen data is on par with baselines with mean AUROC $0.76\pm 0.10$ (all $p>0.05$), while outperforming the NIH baseline ($p=0.009$). Similarly, our method performs comparably to the existing methods.
Table 2: Mean AUROC scores on NIH, CheXpert and MIMIC test sets for non-IID, class-heterogeneous experiment. Our method is compared with others using paired t-tests. Best performing method is bolded.
NIH CheXpert MIMIC
NIH $0.82\pm 0.06^{\text{ns}}$
CheXpert - $0.74\pm 0.07^{*}$
PFL
FedAvg $0.79\pm 0.07^{***}$ $0.73\pm 0.07^{***}$
FedBN $0.81\pm 0.06^{\text{ns}}$ $0.74\pm 0.06^{*}$
FedBN+ $0.81\pm 0.06^{\text{ns}}$ $0.73\pm 0.07^{*}$
Central $0.71\pm 0.13^{**}$ $0.72\pm 0.07^{***}$ $0.65\pm 0.10^{***}$
FL
FedAvg $0.70\pm 0.10^{***}$ $0.68\pm 0.08^{***}$ $0.64\pm 0.09^{***}$
FedBN+ $0.71\pm 0.12^{***}$ $0.68\pm 0.09^{***}$ $0.66\pm 0.10^{***}$
FL + PL
FedAvg $0.77\pm 0.09^{***}$ $0.68\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
FedBN+ $0.78\pm 0.07^{***}$ $0.71\pm 0.07^{***}$ $0.68\pm 0.09^{***}$
FL + SA
FedAvg $0.80\pm 0.06^{*}$ $0.73\pm 0.07^{**}$ $0.70\pm 0.09^{\text{ns}}$
PL: Partial Loss, SA: Surgical Aggregation
(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
[…]
Table 3: Sub-analysis of mean AUROC scores for shared and unique classes of the NIH and CheXpert datasets for non-IID, class-heterogeneous experiment. Our method is compared with others using paired t-tests. Best performing method is bolded.
NIH CheXpert
Shared Classes Unique Classes Shared Classes Unique Classes
NIH $0.84\pm 0.06^{\text{ns}}$ $0.80\pm 0.07^{\text{ns}}$ $0.70\pm 0.08^{**}$
CheXpert $0.81\pm 0.07^{*}$ $0.76\pm 0.07^{\text{ns}}$ $0.71\pm 0.05^{**}$
PFL $0.83\pm 0.06^{**}$ $0.80\pm 0.07^{\text{ns}}$ $0.76\pm 0.07^{\text{ns}}$ $0.70\pm 0.05^{**}$
FL $0.82\pm 0.06^{**}$ $0.60\pm 0.06^{***}$ $0.73\pm 0.09^{**}$ $0.62\pm 0.05^{**}$
FL + PL $0.82\pm 0.05^{*}$ $0.73\pm 0.06^{***}$ $0.74\pm 0.07^{*}$ $0.67\pm 0.04^{**}$
FL + SA $\mathbf{0.84}\pm 0.06$ $\mathbf{0.78}\pm 0.07$ $\mathbf{0.77}\pm 0.09$ $\mathbf{0.75}\pm 0.07$
PL: Partial Loss, SA: Surgical Aggregation
(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
Table 4: Sub-analysis of mean AUROC scores for NIH and CheXpert classes on the MIMIC test for non-IID, class-heterogeneous experiment. Our method is compared with others using paired t-tests. Best performing method is bolded.
MIMIC
NIH Classes CheXpert Classes
NIH $**0.73\pm 0.10^{\text{**}}$
CheXpert - $0.67\pm 0.11^{*}$
PFL
FedAvg $0.74\pm 0.11^{***}$ $0.69\pm 0.10^{\text{ns}}$
FedBN $0.70\pm 0.09^{***}$ $0.67\pm 0.10^{\text{ns}}$
FedBN+ $0.74\pm 0.10^{**}$ $0.68\pm 0.11^{\text{ns}}$
Central $0.74\pm 0.09^{**}$ $0.66\pm 0.11^{**}$
FL
FedAvg $0.72\pm 0.10^{***}$ $0.62\pm 0.06^{**}$
FedBN+ $0.75\pm 0.10^{**}$ $0.63\pm 0.08^{*}$
FL + PL
FedAvg $0.74\pm 0.09^{**}$ $0.66\pm 0.09^{*}$
FedBN+ $0.74\pm 0.10^{**}$ $0.64\pm 0.10^{*}$
FL + SA
FedAvg $0.75\pm 0.09^{\text{ns}}$ $0.68\pm 0.10^{\text{ns}}$
FedBN+ $\mathbf{0.76}\pm 0.10$ $\mathbf{0.70}\pm 0.12$
PL: Partial Loss, SA: Surgical Aggregation
(ns: $p>0.05$, *: $p\leq 0.05$, **: $p\leq 0.01$, ***: $p\leq 0.001$)
On the NIH test set, surgical aggregation performs comparably to the baseline models with mean AUROC $0.81\pm 0.07$ (all $p>0.05$), apart from the FedAvg PFL baseline ($p\leq 0.001$). Furthermore, our method outperforms the existing methods (all $p\leq 0.002$). On the CheXpert test set, surgical aggregation outperforms all baselines (all $p\leq 0.03$) and existing methods (all $p\leq 0.004$) with mean AUROC $0.76\pm 0.08$. Finally, on the external MIMIC dataset, surgical aggregation outperforms all existing methods with mean AUROC $0.71\pm 0.10$ (all $p\leq 0.001$).
[…]
When the mean AUROC scores of the NIH and CheXpert classes is measured separately for the MIMIC test set, we observe that surgical aggregation generalizes better than the baselines and existing methods on NIH classes with mean AUROC $0.76\pm 0.10$ (all $p\leq 0.006$) (4). Similarly, on CheXpert classes, our method yields mean AUROC $0.70\pm 0.12$ and outperforms the existing methods (all $p\leq 0.02$), while performing comparably to the baselines (all $p>0.05$), except the CheXpert baseline ($p=0.04$).
4 Discussion
Class-heterogeneity is a significant challenge in FL, with most FL methods relying on assumptions that all clients either share the same classes, know the classes of other clients in the network, or have access to a fully annotated public dataset. However, there is a significant gap in literature for methods that address class-heterogeneity. Our results show that surgical aggregation can train a global model using distributed, class-heterogeneous datasets.
In the absence of class-heterogeneity, our method performs on par with well-established approaches in literature. However, in the presence of class-heterogeneity, our method continues to maintain high-performance and outperforms the baselines. It also scales across both IID and non-IID settings. On the other hand, existing methods fall short and scale poorly in non-IID settings. For example, while FL with partial loss is a close second in performance with IID data, its performance significantly degrades in non-IID settings. This is due to feature shift as local updates from a client for non-local classes leads to client drift and performance degradation on non-shared classes (Li et al., 2021; Cohen et al., 2020). Unlike our method, FL with partial loss requires clients to account for non-local classes and does not support heterogeneous local classification heads. Therefore, selective aggregation is critical to solve class-heterogeneity in non-IID settings.
Another advantage of our method is its ability to learn a global model that can predict all classes observed in the network. While it is popular approach in non-IID settings, PFL entails $K$ clients learning $K$ personalized models. This has interesting implications in medical imaging: 1) Each personalized model will yield different predictions for class shared between two clients, thus, raising the question of which model’s prediction should be trusted. 2) PFL emphasizes local optimization over generalization, thus making personalized models susceptible to poor generalizability (Zhang et al., 2023).
These advantages enable surgical aggregation to not just facilitate greater collaboration between institutes focusing on different tasks in the same domain, but also enable the countless ”narrowly” focused datasets (e.g., those curated for competitions) to be leveraged to collectively train models with a complete set of abnormalities.
Despite our results demonstrating the advantages of surgical aggregation, our work has certain limitations: 1) There are no established benchmarks for class-heterogeneity beyond external validation. 2) Our method requires clients to share the feature extractor, an important consideration for device heterogeneity. 3) Our experiments were conducted on a simulated FL with a fixed $E=1$ epochs before communication. 4) Our experiments lacked cross-validation and the initial split of data could influence our method’s performance. 5) Due to the primary focus on class-heterogeneity, our ablation study did not evaluate the effect of class-imbalance or non-IID data. For future work, we intend to explore class-heterogeneity using real-world FL setups (e.g., Flower (Beutel et al., 2020)) and further validate our method across different domains and tasks (such as 3D segmentation).
Appendix A Convergence Analysis
[…]
Let us consider that each client $k$ contains private data $\{\mathbf{x}_{k},y_{k}\}$ containing $n_{k}$ samples (indexed by $i$) and is initiated with parameters $\mathbf{w}_{0,k}$ updated using gradient descent.
We primarily focus this analysis on the convergence of selective aggregation to dynamically build the global classification head. In this case, our distributed optimization problem is defined as:
$\min_{\mathbf{w}}\left\{(G\circ F)^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}(G\circ F)_{k}^{*}(\mathbf{w})\right\}$ (A.1)
@@ Appendix A Convergence Analysis
$(G\circ F)_{k}^{*}\triangleq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})$ (A.2)
where $L$ is the loss function (e.g., binary cross-entropy) (Li et al., 2019).
@@
$L(\mathbf{w},\mathbf{x}_{k,i},y_{k,i})=\frac{1}{|C_{k}|}\sum_{c\in C_{k}}\ell(\mathbf{w},\mathbf{x}_{k,i},(y_{k,i})_{c})$ (A.3)
@@
    Optimizing the global feature extractor $F$ by locally optimizing client feature extractors $F_{k}$, such that:
    $\min_{\mathbf{w}}\left\{F^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}^{*}(\mathbf{w})\right\}$ (A.4)
@@
    Optimizing the global classifier $G$ by locally optimizing client classifiers $G_{k}$, such that:
    $\min_{\mathbf{w}}\left\{G^{*}(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}G_{k}^{*}(\mathbf{w})\right\}$ (A.5)
@@
Proof: Like conventional FL approaches, in our approach the feature extractor weights $\mathbf{w}_{t,k}^{(F)}$ are always shared with all clients at the end of each communication round, such that:
$\mathbf{w}_{t}^{(F)}$ $=f(\{\mathbf{w}_{t,k}^{(F)}\ \text{: }\forall k\})$ (A.6)
$\mathbf{w}_{t,k}^{(F)}$ $=\mathbf{w}_{t}^{(F)}\ \forall k$ (A.7)
In this case, the convergence of $\mathbf{w}_{t,k}^{(F)}$ is well-established in prior literature (Li et al., 2019, 2021). However, it is not clear whether class-heterogeneity affects the convergence of $\mathbf{w}_{t,k}^{(F)}$. The optimization of the global sub-objective $F^{*}$ is controlled by the optimization of the local objective $(G\circ F)^{*}_{k}$. Therefore, the optimization of $F^{*}$ is ensured if $F_{k}^{*}\triangleq(G\circ F)^{*}_{k}$ is optimized. While the optimization of $(G\circ F)^{*}_{k}$ is dependent on classes $C_{k}$, it is independent of the optimization of classes $c\in C\backslash C_{k}$ (Lemma 1). In other words, the optimization of $F^{*}$ is driven by the optimization of $F_{k}^{*}$, regardless of any class-heterogeneity.
Since the classifier weights are selectively aggregated, we analyze the convergence of the global classifier (i.e., $G^{*}$) with respect to each individual class $c\in C$.
@@
Proof: Let $K^{\prime}$ be the subset of clients in the network sharing class $c\in C$. In our approach, the class-specific weights $(\mathbf{w}_{t,k}^{(G)})_{c}$ for clients $k\in K^{\prime}$ are shared with the clients at the end of each communication round, such that:
$(\mathbf{w}_{t}^{(G)})_{c}$ $=mean(\{(\mathbf{w}_{t,k}^{(G)})_{c}\ \text{if }c\in C_{k}\ \text{: }\forall k\})$ (A.8)
$(\mathbf{w}_{t,k}^{(G)})_{c}$ $=(\mathbf{w}_{t}^{(G)})_{c}\ \forall k$ (A.9)
@@
Putting together Lemmas 3 and 4, we prove that the optimization of global objective $(G\circ F)^{*}$ with respect to class $c\in C$ is always ensured and is derived from previously established convergence analyses, regardless of class-heterogeneity in the network (Li et al., 2019, 2021). Therefore, using Lemma 1, we can confer:
Theorem 1. The optimization of global objective $(G\circ F)^{*}$ is ensured if local objective $(G\circ F)_{k}^{*}$ is optimized for all classes $c\in C_{k}$ using model aggregation strategy $f$ and the optimization is independent of any class-heterogeneity present in the network.